OpenStudy (anonymous):

The perimeter of the following rectangle is 40 inches. What is the length of w? 2.9 in. 6 in. 12 in. 26 in.

OpenStudy (anonymous):



OpenStudy (ankit042):

perimeter = 2(l+w) 40 = 2(14+w)

OpenStudy (nirmalnema):

6 in